OpenStudy (anonymous):

what property allows the statement 3 = x to be written x = 3? commutative - addition distributive associative - multiplication symmetric commutative - multiplication associative - addition identity - addition

OpenStudy (anonymous):

are thosemultiple choces

OpenStudy (anonymous):

yeah lol

OpenStudy (anonymous):

@rishavraj

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (mathstudent55):

commutative addition: a + b = b + a distributive: a(b + c) = ab + ac associative multiplication: (ab)c = a(bc) symmetric: if a = b, then b = a commutative multiplication: ab = ba associative addition: (a + b) + c = a + (b + c) identity addition: a + 0 = a identity multiplication: a * 1 = a additive inverse: a + (-a) = 0 multiplicative inverse: a * (1/a) = 1 transitive: If a = b, and b = c, then a = c.